FAR-INFRARED LASER MAGNETIC RESONANCE SPECTROSCOPY OF VIBRATIONALLY EXCITED NH

\underline{M. JACKSON}^{a}, L. R. ZINK, Department of Physics, University of Wisconsin-La Crosse, La Crosse, WI 54601; J. FLORES-MIJANGOS, Instituto de Ciencias Nucleares, UNAM, México, D.F.; A. ROBINSON and J. M. BROWN, Physical and Theoretical Chemistry Laboratory, University of Oxford, Oxford, OX1 3QZ, United Kingdom.

The laser magnetic resonance (LMR) spectroscopic technique has been used to investigate the rotational spectrum of the NH radical in its ^{3}Σ^{−} ground electronic state. An optically pumped molecular laser was used as a source of radiation in the 26.7683 to 123.3066 cm^{−1} region of the far-infrared. The molecular transitions were brought into resonance with the laser frequency using magnetic fields up to 2 T. Using twenty-seven laser lines, spectra were recorded in both parallel (ΔM_{J} = 0) and perpendicular (ΔM_{J} = ±1) polarizations. Observation of the N′ ← N′′ = 1 ← 0, 2 ← 1, 3 ← 2 and 4 ← 3 transitions in the first and second excited vibrational levels of NH allows its molecular parameters (particularly the hyperfine constants) to be refined.

aSupport from the National Science Foundation (under Grant No. #0200746) is gratefully acknowledged.